9. The market for pizza is perfectly competitive and has 1,000 firms. Each firm is identical. Describe each firm in long-run12. Bills Bakery has a fire and Bill loses some of his cost data, AFC ATC AVC 100 10 The bits of paper that he recovers afte14. The table shows the total cost of producing sweaters in Henrys factory. Quantity (sweaters per hour) Total cost (dollars
Answer #1

9> A

In the long run, in a perfect competition, firms make zero economic profit.

10> HHI index is the sum of the square of the shares = 0.35^2+0.2^2+0.25^2+0.15^2+0.05^2=0.25

11> In the long run, firms will produces at the lowest ATC, so 200 output per firm with price 10. the quantity demanded is 10000, total number of firms will be 10000/200=50

A is correct

12> Fixed cost is AFCx quantity = 120*10=1200

So, AFC at 20 units will be 1200/20=60 and AVC will be 150-60=90

As HOMEWORKLIB rule, first 4 questions have been answered.